# Anisotropy of Nanohertz Gravitational Wave Background and Individual Sources from Supermassive Binary Black Holes Based on Cosmological Simulation

### v2

@@ Abstract
Several pulsar timing array (PTA) groups have recently claimed the detection of nanohertz gravitational wave background (GWB), but the origin of this gravitational wave (GW) signal remains unclear. Nanohertz GWs generated by supermassive binary black holes (SMBBHs) are one of the most important GW sources in the PTA band. Utilizing data from cosmological simulation, we construct a comprehensive dataset of SMBBHs within a mock observable universe incorporates the cosmic large-scale structure. We carry out an exhaustive analysis of the distribution characteristics of these merger events, as well as the GWB signals they produce. Specifically, we predict the characteristic amplitude of GWB to be $h_{c}=4.85\times 10^{-16}$ at the frequency of ${\rm yr^{-1}}$, while the energy density of GWB signal exhibit an anisotropic part with $C_{1}/C_{0}\approx 2.50\times 10^{-3}\pm 2.04\times 10^{-3}$. We study the clustering pattern of the positional distribution of SMBBHs, and found that they show similar behavior with that of galaxies on relatively small scales. Furthermore, for the upcoming Chinese Pulsar Timing Array (CPTA) and Square Kilometre Array (SKA)-PTA, we predict the spatial distribution, numbers and signal-to-noise ratio (SNR) distribution of individual GW sources that may be detected with SNR$>$8, and study the anisotropic properties in the spatial distribution of these individual GW sources. We finally investigated the impact of lensing effects and found that their influence is rather limited.
@@ 1 Introduction
Precision timing of an ensemble of millisecond pulsars, known as pulsar timing arrays (PTA), offers a unique approach to detect such low-frequency GW signals including GWB and individual sources (Sazhin, 1978; Detweiler, 1979; Blandford et al., 1984; Foster & Backer, 1990; Maggiore, 2008; Sesana et al., 2009; Sesana & Vecchio, 2010; Creighton & Anderson, 2011; Lee et al., 2011; Manchester, 2013; Sesana, 2013a; van Haasteren, 2014; Zhu et al., 2014; Blair et al., 2015; Mingarelli, 2015; Schutz & Ma, 2016; Taylor et al., 2019). Currently, a diverse array of PTA experiments is operational, including the Parkes PTA (PPTA [^1] http://www.atnf.csiro.au/research/pulsar/ppta/; Manchester 2013), the European PTA (EPTA [^2] http://www.epta.eu.org/; Kramer & Champion 2013), the North American Nanohertz Observatory for Gravitational Waves (NANOGrav [^3] http://nanograv.org/; McLaughlin 2013; Ransom et al. 2019), the Indian Pulsar Timing Array (InPTA; Joshi et al. 2018), the Chinese PTA (CPTA; Nan et al. 2011; Smits et al. 2009), and the MeerKAT PTA (MPTA; Miles et al. 2023). Among which, the first three PTA groups have amassed data spanning over a decade in pursuit of potential low-frequency GW signals. They have also collaborated to form the International PTA (IPTA [^4] http://www.ipta4gw.org/; Manchester & IPTA 2013; Brazier et al. 2016; Perera et al. 2019) by sharing data. Looking forward, the Square Kilometer Array (SKA) (e.g., Lazio, 2013; Wang & Mohanty, 2017) is anticipated to discover numerous stable millisecond pulsars (MSPs) and establish a high-sensitivity PTA, henceforth referred to as SKA-PTA, to detect low-frequency GW sources. Recently, CPTA, NANOGrav, EPTA (+InPTA), and PPTA have presented evidence for the presence of a nanohertz GWB with a confidence level of approximately 2-4 $\sigma$(Xu et al., 2023; Agazie et al., 2023a; Reardon et al., 2023; EPTA Collaboration et al., 2023)(see also Arzoumanian et al. 2020). Although the origin of this GWB is still unclear, it is likely that this signal can be attributed to, or at least partly attributed to GWs generated from cosmic inspiralling SMBBHs, indicating that the detection of individual SMBBHs may also soon become a reality (e.g., Chen et al., 2023; Agazie et al., 2023b; Antoniadis et al., 2023).
@@
In recent years several studies have also presented predictions on this GWB based on phenomenological models or simulations (Jaffe & Backer, 2003; Sesana, 2013b; Sesana et al., 2016; Kelley et al., 2017; Yang et al., 2019). The predicted GW strain amplitude ($A_{\rm yr^{-1}}$) lies approximately within the range of $1\times 10^{-16}$ to $5\times 10^{-15}$. The practical GW strain amplitude $A_{\rm yr^{-1}}$) from GWB detected by NANOGRav, CPTA, EPTA and PPTA is around $2\times 10^{-15}$. The predictions regarding the characteristic amplitude of the GWB focus on its isotropic nature, which represents an ideal scenario when the sources of the GWB are extremely numerous that their distribution are nearly continuous, and resulting in the predicted characteristic Hellings and Downs curve that could be observed in PTA experiments (Hellings & Downs, 1983). However, in reality, the number of SMBBHs is finite, leading to a GWB signal comprising both a dominant isotropic component and also anisotropic contributions. The impact of anisotropies on PTA measurements and the extension of the Hellings and Downs curve method to analyze anisotropies in the GWB have been explored in various studies, including Cornish & Sesana (2013); Mingarelli et al. (2013); Taylor & Gair (2013); Cornish & van Haasteren (2014); Mingarelli (2015); Hotinli et al. (2019); Sato-Polito & Kamionkowski (2024). Several PTAs have also provided constraints on the level of anisotropy in the GWB (Taylor et al., 2015).
@@
Although there has been some work exploring the anisotropy property of the GWB signal (Yang et al., 2019; Sato-Polito & Kamionkowski, 2024), they usually rely on randomly distributed SMBBHs without considering the possible effects of the large-scale structure of the universe. The method we use allow us to estimate the anisotropy of GWB in a simulation environment where the cosmic large scale structure is incorporated, thus more reflective to the real universe and offering a more reliable result. In this paper, we intend to employ a semi-analytic galaxy formation model (SAM) grounded on dark matter (DM) halo merger trees derived from the Millennium simulation (Springel et al., 2005; Guo et al., 2011). We will focus on the MR model, specifically the variants proposed by Guo et al. (2011, 2013) (see also Henriques et al. 2015), to construct a comprehensive dataset of SMBBHs within a mock observable universe, from which we forecast the formation rate of SMBBHs, their statistical properties regarding redshifts, mass and mass ratios, their spatial and angular distributions, and the characteristic strain amplitude of the resulting GWB, along with its anisotropic characteristics. We conduct a detailed study on the three dimensional spatial and two dimensional angular distributions of GW sources with the SMBBHs that we constructed in the observable universe, and present the resulting angular power spectrum coefficients and angular two-point correlation functions. We also predict resolvable individual GW sources for future CPTA and SKA-PTA, and show their statistical properties. In addition, we will also make an estimation on how gravitational lensing effect will affect all of our calculational results.
@@
This paper is organized as follows. In Section 2, we introduce the galaxy evolution model and mock data we used, as well as the basic setups and methods for constructing SMBBHs in the observable universe. In Section 3, we present our basic results concerning the isotropic and anisotropic characteristics of the GWB signal, which we derived from the SMBBH data constructed, and also the clustering properties of the GW sources. We further discuss the detection capabilities of different PTAs for individual bright sources in Section 4. In Section 5, we take into account the effects of gravitational lensing and discuss their impact on all of our calculational results derived above. Finally, Section 6 is devoted to conlusions and discussions. Throughout this paper, we adopt that dimensionless Hubble constant $h_{0}=0.704$.
@@ 2 Models
In this section, we introduce how we construct the light cone SMBBH data in the observable universe. The flowchart of the entire process to deal with simulation data and generate simulated GW sources is shown in Figure 1. We will introduce our method related to each part of the flowchart in detail in the following subsections.
Figure 1: A flowchart for constructing GW events in the observable universe from Guo2013a MR mock data.
@@ 2.1 Mock Data
We utilize the results of L-Galaxies semi-analytical galaxy formation runs using the model described in Guo et al. (2011, 2013) stored in the Millennium simulation database Guo2013a. This is a semi-analytical modelling applied to the DM halo merger trees of the Millennium simulations, which follows the formation, evolution and mergers of galaxies in a box with side length of 500 Mpc$/h_{0}$ from $z>14$ to $z=0$. Since SMBH and galactic nuclear activity is closely related to galaxy formation, SAM always encompass the growth and evolution of SMBHs in their modeling. In Guo2013a, “quasar” mode and “radio” mode are considered to grow SMBHs from initial seeds, leading to a co-evolution of SMBHs and their host galaxies that is compatible with observations. We assume that SMBHs will form inspiring binaris following the merger of their host galaxies, so according to the merger history of DM halos and galaxies from Guo2013a simulation result, we can judge when and where a SMBBH will form. In this way, we get the SMBBH merger events distributed in one simulation box, each event carries information including their locations in the simulation box, redshifts, host galaxies, progenitor black hole masses and descendant black hole mass.
2.2 Observable Universe
@@
The mock simulation data are output in 61 snapshots at different redshifts in a box with side length 500 Mpc$/h_{0}$. However, instead of SMBBHs at specific redshifts in a box with limited size, signals observed by PTA originate from SMBBHs distributed along the light cone, with larger distance to us at higher redshifts. These SMBBHs constitutes the observable universe centered at the earth, and the distance of them may exceed the side length of the simulation box at relatively high redshifts. In order to construct this observable universe with mock data, and considering the universe is uniform and isotropic on large scale ($>100$ Mpc), we adopt periodic boundary conditions, i.e., we piece together the simulation box in a repeating pattern to generate a universe that can theoretically stretch out indefinitely in all three directions. We set the position of the observer to be at the origin, which is a vertex of the simulation box with coordinates (0,0,0), e.g., see the blue sphere in Figure 2 and Figure 7. For each redshift snapshot, we calculate the corresponding comoving distance. We then draw a spherical shell at this distance with certain thickness, which we chose to be the comoving distance difference between neighboring redshifts. Any SMBBH event that falls within the shell corresponds to its redshift is considered to be within the light cone and are saved as observable GW source events. In Figure 2, we illustrate this process for the initial three redshift snapshots. A gentle reminder is that we fully utilized the total 61 redshift snapshots in our analysis instead of the initial three as depicted in the figure.
Figure 2: An illustration of constructing the observable universe using periodic boundary conditions for the initial three redshift snapshots. The upper three plots represent the periodic simulation boxes of the initial three redshift snapshots respectively. The spherical shells representing the extent of the light cone are shown in each plot. SMBBHs that fall within these shells are selected as light cone events. By collecting the light cone events from different redshifts, the observable universe is constructed.
@@ 2.3 GW Sources
We ultimately obtained $N_{\rm total}\approx 3.027\times 10^{7}$ SMBBH events in the observable universe with methods described in the previoues subsection. We consider these SMBBHs to serve as GW sources whose signals could be detectable in the PTA frequency band. To predict the characteristics of the GWB signal from these GW sources, we still need to determine the GW strain and frequency for each SMBBH events, the method is described below.
2.3.1 GW Strain
Given that the pulsars utilized by PTA may be situated at various locations with diverse inclinations, the orientation of the PTA itself can be considered arbitrary. Furthermore, the SMBBHs under observation can also exhibit a wide range of inclinations. Hence we adopt the following formula to estimate the average GW strain for each SMBBH:
@@ 2.3.1 GW Strain
where $d_{\rm L}$ is the luminosity distance of GW source, $\mathcal{M}_{z}=\mathcal{M}(1+z)$ is the redshifted mass, with chirp mass $\mathcal{M}=\frac{(m_{1}m_{2})^{3/5}}{M^{1/5}}$, and $m_{1}$, $m_{2}$ are the masses of the two components of this SMBBH, $M=m_{1}+m_{2}$ is the total mass of this SMBBH. Observational frequency $f_{\rm obs}=\frac{f_{\rm r}}{1+z}$, where $f_{\rm r}$ is the GW frequency in the rest frame of GW source. Here we assume the orbit of each SMBBH is circular for simplicity. As for how the GW frequency is generated is described in the next subsubsection.
2.3.2 Frequency Generating
The PTA sensitive frequency range is limited by two parameters: observational duration $T_{\rm obs}$ and observational cadence $\Delta t$. Under these two parameters the detectable frequency range can be determined as $\frac{1}{T_{\rm obs}}<f_{\rm obs}<\frac{1}{\Delta t}$. When considering an observational duration of a few decades, the lower limit of the frequency band is approximately $10^{-9}$z. The frequencies of the GW signals from individual sources are determined by a variety of factors, including the orbital separation of the two black holes, their individual masses, the redshift of the source, and so on. Mock data from SAM do not contain these information, so we need to assign frequencies to each SMBBH system randomly. We adopt the quadrupole approximation, accordding to which the residual time within a unit logarithmic frequency interval is given by (Maggiore, 2008)
$\frac{dt_{\rm r}}{d\ln f_{\rm r}}=\frac{5}{96}\pi^{-8/3}\left(\frac{G\mathcal{M}}{c^{3}}\right)^{-5/3}f_{\rm r}^{-8/3}.$ (2)
We assume that the probability of a SMBBH system being at a certain orbital frequency is directly proportional to the residual time, thus the probability distribution of the rest frame GW frequency from BBH merger is
$P(\ln f_{\rm r})=\frac{dP}{d\ln f_{\rm r}}\propto\mathcal{M}^{-5/3}f_{\rm r}^{-8/3},$ (3)
or
$\frac{dP}{df_{\rm r}}\propto\mathcal{M}^{-5/3}f_{\rm r}^{-11/3}.$ (4)
We restrict the distribution of $f_{\rm r}$ to be between $f_{0}$ and $f_{\rm ISCO}$, where $f_{0}=10^{-9}$ Hz is roughly the lower limit of PTA frequency band and $f_{\rm ISCO}$ is the GW frequency when the SMBBH system is at the inner-most stable circular orbit. $f_{\rm ISCO}$ may vary among different SMBBH systems. We calculate $f_{\rm ISCO}$ as follows: According to Kepler’s third law, $GM=a^{3}\omega_{\rm orb}^{2}=\pi^{2}a^{3}f_{\rm r}^{2}$ with $M$ the total mass, for spinless SMBBH, the radius of the inner-most stable circular orbit $a=\frac{6GM}{c^{2}}$, thus GW frequency from this orbit is given by (Maggiore, 2008; Sathyaprakash & Schutz, 2009):
$f_{\rm ISCO}=\frac{c^{3}}{\pi 6^{\frac{3}{2}}GM}\approx 4.4\times 10^{-7}{\rm Hz}\left(\frac{10^{10}M_{\odot}}{M}\right).$ (5)
According to the probability distribution of Eq. (4), we randomly generate frequency for each of the GW sources (see also Figure 5) with Monte Carlo method.

### v3

# Anisotropy of Nanohertz Gravitational Wave Background and Source Clustering from Supermassive Binary Black Holes Based on Cosmological Simulation

@@ Abstract
Several pulsar timing array (PTA) groups have recently claimed the detection of nanohertz gravitational wave background (GWB), but the origin of this gravitational wave (GW) signal remains unclear. Nanohertz GWs generated by supermassive binary black holes (SMBBHs) are one of the most important GW sources in the PTA band. Utilizing data from cosmological simulation, we generate multiple realizations of a mock observable universe that self-consistently incorporates the cosmic large-scale structure, enabling a robust statistical analysis of SMBBH populations and their GW signatures. We systematically investigate the merger event distributions and both the isotropic and anisotropic properties of the resulting GWB signals under different hardening timescales. Specifically, we calculate the characteristic amplitude of the GWB signal, and the angular power spectrum for both the total energy density and energy density in different frequency bins accounting for cosmic variance through different realizations. We also study the clustering pattern of the positional distribution of SMBBHs to examine whether they can reproduce the large-scale structure of galaxies. Furthermore, for the upcoming Chinese Pulsar Timing Array (CPTA) and Square Kilometre Array (SKA)-PTA, we predict the numbers and signal-to-noise ratio (SNR) distributions of resolvable individual GW sources that may be detected with SNR$>$8. We finally investigated the impact of weak lensing effects and found that their influence on the basic characteristics of the GWB signal and individual sources is rather limited.
@@ 1 Introduction
Precision timing of an ensemble of millisecond pulsars, known as pulsar timing arrays (PTA), offers a unique approach to detect such low-frequency GW signals including GWB and individual sources (Sazhin, 1978; Detweiler, 1979; Blandford et al., 1984; Foster & Backer, 1990; Maggiore, 2008; Sesana et al., 2009; Sesana & Vecchio, 2010; Creighton & Anderson, 2011; Lee et al., 2011; Manchester, 2013; Sesana, 2013a; van Haasteren, 2014; Zhu et al., 2014; Blair et al., 2015; Mingarelli, 2015; Schutz & Ma, 2016; Taylor et al., 2019). Currently, a diverse array of PTA experiments is operational, including the Parkes PTA (PPTA [^1] http://www.atnf.csiro.au/research/pulsar/ppta/; Manchester 2013; Reardon et al. 2023), the European PTA (EPTA [^2] http://www.epta.eu.org/; Kramer & Champion 2013; EPTA Collaboration et al. 2023), the North American Nanohertz Observatory for Gravitational Waves (NANOGrav [^3] http://nanograv.org/; McLaughlin 2013; Ransom et al. 2019; Agazie et al. 2023a), the Indian Pulsar Timing Array (InPTA; Joshi et al. 2018; EPTA Collaboration et al. 2023), the Chinese PTA (CPTA; Nan et al. 2011; Smits et al. 2009; Xu et al. 2023), and the MeerKAT PTA (MPTA; Miles et al. 2023). Among which, the first three PTA groups have amassed data spanning over a decade in pursuit of potential low-frequency GW signals. They have also collaborated to form the International PTA (IPTA [^4] http://www.ipta4gw.org/; Manchester & IPTA 2013; Brazier et al. 2016; Perera et al. 2019) by sharing data. Looking forward, the Square Kilometer Array (SKA) (e.g., Lazio, 2013; Wang & Mohanty, 2017) is anticipated to discover numerous stable millisecond pulsars (MSPs) and establish a high-sensitivity PTA, henceforth referred to as SKA-PTA, to detect low-frequency GW sources. Recently, CPTA, NANOGrav, EPTA (+InPTA), and PPTA have presented evidence for the presence of a nanohertz GWB with a confidence level of approximately 2-4 $\sigma$(Xu et al., 2023; Agazie et al., 2023a; Reardon et al., 2023; EPTA Collaboration et al., 2023)(see also Arzoumanian et al. 2020). Although the origin of this GWB is still unclear, it is likely that this signal can be attributed to, or at least partly attributed to GWs generated from cosmic inspiralling SMBBHs, indicating that the detection of individual SMBBHs may also soon become a reality (e.g., Chen et al., 2023; Agazie et al., 2023b; Antoniadis et al., 2023).
@@
In recent years several studies have also presented predictions on this GWB based on phenomenological models or simulations (Jaffe & Backer, 2003; Sesana et al., 2008; Sesana, 2013b; Sesana et al., 2016; Kelley et al., 2017; Yang et al., 2019; Gardiner et al., 2024; Allen et al., 2024; Sah et al., 2024; Semenzato et al., 2024; Grimm et al., 2025). The predicted GW strain amplitude ($A_{\rm yr^{-1}}$) lies approximately within the range of $1\times 10^{-16}$ to $5\times 10^{-15}$. The practical GW strain amplitude $A_{\rm yr^{-1}}$ from GWB detected by NANOGRav, CPTA, EPTA and PPTA is around $2\times 10^{-15}$. The predictions regarding the characteristic amplitude of the GWB focus on its isotropic nature, which corresponds to an idealized case where the GWB sources are sufficiently abundant to approximate a continuous spatial distribution, and thus results in the characteristic Hellings and Downs curve that could be observed in PTA experiments (Hellings & Downs, 1983). However, in reality, the number of SMBBHs is finite, leading to a GWB signal comprising both a dominant isotropic component and also anisotropic contributions (Allen et al., 2024). The impact of anisotropies on PTA measurements and the extension of the Hellings and Downs curve method to analyze anisotropies in the GWB have been explored in various studies, including Cornish & Sesana (2013); Mingarelli et al. (2013); Taylor & Gair (2013); Cornish & van Haasteren (2014); Mingarelli (2015); Hotinli et al. (2019); Sato-Polito & Kamionkowski (2024). Several PTAs have also provided constraints on the level of anisotropy in the GWB (Taylor et al., 2015).
@@
There has been some work exploring the anisotropy property of the GWB signal by utilizing random Poisson sampling of GW sources (Yang et al., 2019; Sato-Polito & Kamionkowski, 2024; Gardiner et al., 2024); by using cosmological simulations to estimate the annisotropic characteristics of GWB signal or its cross correlation with galaxies (Sah et al., 2024; Semenzato et al., 2024); or by using theoretical framework to quantify the impact of GW source discreteness and clusterings on the Hellings and Downs curve (Allen et al., 2024; Grimm et al., 2025). In this work, we intend to employ a semi-analytic galaxy formation model (SAM) grounded on dark matter (DM) halo merger trees derived from the Millennium simulation (Springel et al., 2005; Guo et al., 2011). We will focus on the SAM variants proposed by Guo et al. (2011, 2013) (see also Henriques et al. 2015) to construct multiple realizations of a comprehensive dataset of SMBBHs within the mock observable universe, from which we forecast the statistical properties of SMBBHs regarding their redshifts, mass and mass ratios, their spatial and angular distributions, and the characteristic strain amplitude of the resulting GWB, along with its anisotropic characteristics. Our method allows us to incorporate the cosmic large-scale structure and hardening timescales in a self-consistent manner at the level of light-cone SMBBH events construction. In addition, we study the two dimensional angular distributions of GW sources with the SMBBHs that we constructed in the observable universe, and present the resulting angular power spectrum coefficients and angular two-point correlation functions. We also predict resolvable individual GW sources for future CPTA and SKA-PTA, and show their statistical properties. Finally, we will also make an estimation on how gravitational lensing effect will affect all of our calculational results.
@@
This paper is organized as follows. In Section 2, we introduce the galaxy evolution model and mock data we used, as well as the basic setups and methods for constructing light-cone SMBBHs in the observable universe. In Section 3, we present our basic results concerning the isotropic and anisotropic characteristics of the GWB signal, which we derived from the SMBBH data constructed, and also the clustering properties of the GW sources. We further discuss the detection capabilities of different PTAs for individual bright sources in Section 3.4. In Section 4, we take into account the effects of gravitational lensing and discuss their impact on all of our calculational results derived above. Finally, Section 5 is devoted to conlusions and discussions. Throughout this paper, we adopt that dimensionless Hubble constant $h_{0}=0.704$.
@@ 2 Models
In this section, we introduce how we construct the light-cone SMBBH data in the observable universe. The flowchart and schematic of the entire process to deal with simulation data and generate light-cone GW sources are shown in Figure 1. We will introduce our method related to each part of the flowchart in detail in the following subsections.
Figure 1: Left panel: a flowchart for the entire process to deal with simulation data and construct light-cone GW events in the observable universe from Guo2013a mock galaxy data based on Millennium Run; Right panel: a schematic diagram illustrating the method for determining whether a SMBBH event lies within the light cone. Here, $t_{0}$, $t_{1}$ and $t_{2}$ are the cosmic times when the two progenitor galaxies merge, the SMBBH inside them entering PTA band and SMBBH coalescence respectively; $z_{0}$, $z_{1}$ and $z_{2}$ are the corresponding redshifts of these events; $d_{c}(z)$ is the comoving distance related to the redshift $z$. We apply this selection process to each SMBBH in every merger-formed galaxy to obtain all light-cone events.
@@ 2.1 Mock Data
We utilize the results of L-Galaxies semi-analytical galaxy formation runs using the model described in Guo et al. (2011, 2013) stored in the Millennium simulation database Guo2013a. This is a semi-analytical modelling applied to the DM halo merger trees of the Millennium simulations, which follows the formation, evolution and mergers of galaxies in a box with side length of 500 Mpc$/h_{0}$ from $z>14$ to $z=0$. The mock simulation data are output in 61 snapshots at different redshifts. From the output data, we can extract those galaxies formed by mergering in each redshift snapshot, each of these galaxies carries information about its location and its two progenitor galaxies. Thus the merger history of galaxies across the entire redshift range can be constructed. Since SMBH and galactic nuclear activity are closely related to galaxy formation, SAM always encompass the growth and evolution of SMBHs in their modeling. In Guo2013a, “quasar” mode and “radio” mode are considered to grow SMBHs from initial seeds, leading to a co-evolution of SMBHs and their host galaxies that is compatible with observations. We assume that SMBHs will form inspiralling binaries following the merger of their host galaxies, so that the merger history of supermassive black holes can be correspondingly constructed following the merger history of galaxies.
2.2 Observable Universe with different hardening timescale
We assume that supermassive black holes will reach the PTA frequency band ($10^{-9}{\rm Hz}-10^{-6}{\rm Hz}$) after undergoing a uniform hardening time following galaxy mergers. We consider three choices for the hardening timescale $\tau_{H}$, i.e., $\tau_{H}=0.1\,\mathrm{Gyr}$, $5\,\mathrm{Gyr}$, and $10\,\mathrm{Gyr}$. We randomly place Earth in 50 different galaxies, as is shown in Figure 2 (displacements at the order of $10$ kpc relative to the galaxy centers are randomly assigned to each Earth position to ensure they are not too close to any SMBBH merger events).
Figure 2: Earth’s positions are randomly distributed across 50 different galaxies, allowing us to generate 50 distinct realizations of the light-cone SMBBH events for each of the three hardening timescales.
@@ 2.2 Observable Universe with different hardening timescale
This allows us to generate 50 different realizations of the SMBBH merger events for each hardening timescale. However, instead of SMBBHs at specific redshifts in a box with limited size, signals observed by PTA originate from SMBBHs distributed along the light cone, with larger distance to us at higher redshifts. These SMBBHs constitutes the observable universe centered at the earth, and the distance of them may exceed the side length of the simulation box at relatively high redshifts. In order to construct this observable universe with mock data, and considering the universe is uniform and isotropic on large scale ($>100$ Mpc), we adopt periodic boundary conditions, i.e., we piece together the simulation box in a repeating pattern to generate a universe that can theoretically stretch out indefinitely in all three directions. For each simulation box and every merger-formed galaxy recorded at certain redshift snapshots, we randomly assign to them a cosmic time between the current redshift snapshot and the previous one to represent the merger time of the two progenitor galaxies. The time when the SMBBH enters the PTA band is then determined by adding an extra hardening time to this merger time. For each SMBBH event, we calculate the evolution time spanning from its entry into the PTA band to final coalescence (i.e., from $(1+z)\times 10^{-9}\rm Hz$ to coalescence, where $z$ is redshift of the SMBBH when it enters the PTA band, and can be determined by converting cosmic time to redshift) according to the following equation:
$\tau=\frac{5}{256}(\pi f_{r})^{-\frac{8}{3}}\left(\frac{GM_{c}}{c^{3}}\right)^{-\frac{5}{3}}$ (1)
This evolutionary timescale spans a specific comoving distance interval, any SMBBH event whose distance to Earth falls within this interval is considered to be within the light cone and are saved as a potential observable GW source event. The distances of these GW events to Earth can determine a specific look-back time. By calculating how much time remains from this look-back time to the time of coalescence, the rest frame frequency of each binary black hole event can be determined by solving for $f_{r}$ with a certain $\tau$ through Equation (1). Here we have assumed the orbits of SMBBHs are circular for simplicity. In the right panel of Figure 1, we also present a schematic diagram of our method for constructing the light-cone events.
@@ 2.3 GW Sources
With methods described in the previous subsection, we ultimately obtained $N_{\rm total}\approx 1\times 10^{8},6\times 10^{6},1\times 10^{4}$ SMBBH events in the observable universe for hardening time scales of $0.1,5,10\,\mathrm{Gyr}$ respectively. Among them, we consider the SMBBHs whose $f_{\rm obs}$ falls between $f_{min}=1/T_{\rm obs}$ and $f_{max}=1/\Delta t$ to serve as GW sources whose signals are detectable for the PTA configurations we will consider in subsequent sections, where $T_{\rm obs}$ is the total observation duration and $\Delta t$ is the observation cadence. Each binary system possesses a unique combination of spatial coordinates, redshift, black hole masses, and are emitting GWs at a certain frequency. We can determine the correponding GW strain amplitude for each SMBBH event with all this information. Given that the pulsars utilized by PTA may be situated at various locations with diverse inclinations, the orientation of the PTA itself can be considered arbitrary. Furthermore, the SMBBHs under observation can also exhibit a wide range of inclinations. Hence we adopt the following formula to estimate the average GW strain for each SMBBH:
@@
where $d_{\rm L}$ is the luminosity distance of GW source, $\mathcal{M}_{z}=\mathcal{M}(1+z)$ is the redshifted mass, with chirp mass $\mathcal{M}=\frac{(m_{1}m_{2})^{3/5}}{M^{1/5}}$, and $m_{1}$, $m_{2}$ are the masses of the two components of this SMBBH, $M=m_{1}+m_{2}$ is the total mass of this SMBBH. Observational frequency $f_{\rm obs}=\frac{f_{\rm r}}{1+z}$, where $f_{\rm r}$ is the GW frequency in the rest frame of GW source.
It should be noted that, under conditions of extended hardening timescales, a SMBBH within a host galaxy may experience multiple subsequent merger processes before coalescence; we did not model the dynamical evolution of such multi-body systems. Instead, we adopt a simplified treatment assuming the new incoming black hole will form a new binary system with the remnant of the prior coalescence. The treatment of sequential mergers as independent GW sources, rather than interconnected dynamical systems, represents a limitation in capturing full self-consistency; however, it provides a tractable framework with enough computational efficiency. Due to the same reason, we also didn’t account for black hole mass evolution after the merger of galaxies. We evaluate the potential impact of this simplification in Appendix A for $\tau_{H}=10\,\mathrm{Gyr}$, in which case mass evolution has the greatest impact.